# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v2

@@ 7 Summary
SPARSE is a simple and efficient Python algorithm for the calculation of amplitudes and resonances for inelastic scattering in nonrelativistic quantum mechanics. SPARSE solves a system of coupled radial Schrödinger equations that represents a partial wave of a two-body scattering process in systems where the particles can change their state as a result of the collision. The SPARSE algorithm is optimized for the application to systems with tens of coupled Schrödinger equations. It reduces the system of differential equations to a nonhomogeneous linear system by approximating the second derivative using the finite-difference method and imposing Dirichlet boundary conditions. It calculates the numerical wavefunctions by solving the system, then compares them to the analytical wavefunctions of scattering states to obtain the $K$-matrix elements. Finally, SPARSE calculates scattering amplitudes and poles from the $K$-matrix elements for real energies.

### v3

@@ 7 Summary
SPARSE is a simple and efficient Python algorithm for the calculation of amplitudes and resonances for elastic or inelastic scattering of particles with spin in nonrelativistic quantum mechanics. SPARSE solves a system of coupled radial Schrödinger equations that represents a partial wave of a two-body scattering process in systems where the particles can change their state as a result of the collision. The SPARSE algorithm is optimized for the application to systems with dozens of coupled Schrödinger equations. It reduces the system of differential equations to a nonhomogeneous linear system by approximating the second derivative using the finite-difference method and imposing Dirichlet boundary conditions. It calculates the numerical wavefunctions by solving the system, then compares them to the analytical wavefunctions of scattering states to obtain the $K$-matrix elements. Finally, SPARSE calculates scattering amplitudes and poles from the $K$-matrix elements for real energies.